OpenStudy (anonymous):

What is the d value of the following arithmetic sequence? 5, 2, -1, -4, -7,…

OpenStudy (anonymous):

-3 because 5-3=2 and 2-3=-1 and -1-3=-4 and so on.

OpenStudy (goformit100):

-3 is the d

OpenStudy (anonymous):

how did yu get that